Circular No. 8208

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 S1 (NEAT) M. Hicks, Jet Propulsion Laboratory (JPL), reports the discovery of a

comet on CCD images obtained via the NEAT project with the Haleakala 1.2-m telescope (discovery observation below), noting the object to be slightly diffuse with a slight elongation to the west on images taken on Sept. 23.6 and 24.5 UT. Following posting on the NEO Confirmation Page, P. Birthwhistle, Great Shefford, U.K., reported that his CCD images taken with a 0.30-m reflector on Sept. 24.0 show a coma with diameter ∼ 10′′ offset toward the southwest with a possible 30′′ tail in p.a. ≈ 250^{o}. J. Mahony, Lafayette, IN, communicates that his images with a 0.30-m Schmidt-Cassegrain reflector on Sept. 24.2 show a diffuse coma and a tail ∼ 10′′ long in p.a. 250^{o}. S. D. Gillam, JPL, obtained confirming images with the 0.6-m telescope at Table Mountain on Sept. 24.2, showing a coma of diameter ≈ 10′′ with mag R = 18.3. Images taken with the 1.06-m KLENOT telescope at Kleť by M. Tichý on Sept. 24.9 show a tail in p.a. 260^{o}.

2003 UT  α_{2000}  δ_{2000}  Mag.
Sept.23.60194  0^{h}35^{m}19.^{s}01  +11^{o}37′32′′.4  18.8

The available astrometry (including LINEAR prediscovery observations from Sept. 4.3 and 20.3), the following elliptical orbital elements, and an ephemeris appear on MPEC 2003-S59.

T = 2004 Mar. 25.984 TT  ω = 175.^{o}232
e = 0.42790  Ω = 241.130 }2000.0
q = 2.60068 AU  i = 5.937
a = 4.54586 AU n^{o} = 0.101690 P = 9.69 years

SUPERNOVA 2003ii IN MCG +06-6-53 Further to IAUC 8205, T. Boles reports the discovery of an apparent

supernova (mag 18.4) on unfiltered CCD images taken on Sept. 23.030 and 23.922 UT. The new object is located at α = 2^{h}33^{m}04^{s}.80, δ = +35^{o}30′29′′.7 (equinox 2000.0), which is ≈ 8′′.1 east and 2′′.1 north of the center of MCG +06-6-53. SN 2003ii is not present on Boles’ images from Aug. 14 and Sept. 2 (limiting mag 19.5), and it is not present on Palomar Sky Survey red (1986) or blue (1987) plates.

2003 September 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green